3-26 The Blacksburg Company is separating materials handling cost into fixed....

5. E (c) Identify the statistical factors which are used to test a regression equation for good- ness of fit and, for each item identified, indicate whether a high or low value describes a good fit (d) Explain what each of the following terms means with respect to regression analysis. 1. Linearity within a relevant range 2. Constant variance (homoscedasticity) 3. Serial correlation 4. Normality 5. Multicollinearity (IMA adapted) High-Low Cost Estimation and Simple Regression Analysis The components. The following data have been collected over a five-month peri amounts are in thousands) 3-26 Blacksburg Company is separating materials handling cost into fixed and variable Direct Cost of Materials Month Labor Hours Materials Issued Handling Costs $1.60 1.70 1.80 1.90 2.00 $7 Required (1) Determine the fixed and variable materials handling cost by the high-low method using direct labor hours as the measure of volume (2) Determine the fixed and variable materials handling cost by the method of least squares using direct labor hours as the measure of volume (3) Repeat (2) using cost of materials issued as the measure of volume (4) Which of the two measures of volume in (2) and (3) seems preferable? Why?

If possible, can you please explain without Microsoft Excel. Thank you very much.
Answer #1

(1): b (variable cost) = (y2-y1)/(x2-x1).

Y2 = total cost at the highest level of activity, y1 = total cost at the lowest level of activity, x2 = highest level of direct labor hours and x1 = lowest level of direct labor hours.

Thus b = (2-1.6)/(8-4)

= $0.10

Fixed cost = y2-bx2 = 2 – (0.1*8) = 1.2.

Thus cost formula is 1.2+0.10x. Hence fixed costs is $1.2 and variable cost is $0.10 per direct labor hour

(2):

x (direct labor hours) y (material handling cost) xy x^2
4 1.60 6.40 16.00
5 1.70 8.50 25.00
6 1.80 10.80 36.00
7 1.90 13.30 49.00
8 2.00 16.00 64.00
Total 30.00 9.00 55.00 190.00

b = n∑xy – (∑x)(∑y)/ n∑x^2 - (∑x)^2

= 5*55 - (30*9)/5*190 - (30^2)

= 0.10

a = ∑y - b∑x/n

= (9-0.1*30)/5

= 1.2.

Thus variable cost = $0.10 per direct labor hour and fixed cost = $1.2

(3) Here x = cost of material issued

x (cost of material issued) y (material handling cost) xy x^2
7 1.60 11.20 49.00
5 1.70 8.50 25.00
4 1.80 7.20 16.00
8 1.90 15.20 64.00
6 2.00 12.00 36.00
Total 30.00 9.00 54.10 190.00

b = n∑xy – (∑x)(∑y)/ n∑x^2 - (∑x)^2

= 5*54.10 - 30*9/5*190 - (30^2)

= 0.01

a = ∑y - b∑x/n

= (9-0.01*30)/5

= 1.74

Thus variable cost = $0.01 per cost of material issued and fixed cost is $1.74

(4): The measure of volume in (2) will be preferred as material handling costs will be more dependent on direct labor hours than cost of materials issued. Direct labor hours will be a better cost driver as material handling costs vary proportionately with quantity of materials. The higher the quantity of materials higher will be the direct labor hours required.